OpenStudy (anonymous):

Please tell me how to do this problem Evaluate : sin(1/4 pie) A) 1 B) 0 C) sqrt 2 /2 D None of these

OpenStudy (anonymous):

C) sqrt 2 /2

OpenStudy (anonymous):

How did u get this?

OpenStudy (anonymous):

look at the unit circle on the last page of this cheat sheet.

OpenStudy (anonymous):

c

OpenStudy (anonymous):

you will see that at \[\frac{\pi}{4}\] both the first and second coordinate on the unit circle is \[\frac{\sqrt{2}}{2}\] so both sine and cosine are the same there

OpenStudy (anonymous):

you can also do it directly using a triangle but this one comes up so often it is best to memorize it

OpenStudy (anonymous):

ok then how would i do a problem like tan( -337 pie /4 )

OpenStudy (anonymous):

oh lord. first of all you best find a point between 0 and 2 pi that is coterminal with \[\frac{-337\pi}{4}\]

myininaya (myininaya):

pie?

OpenStudy (anonymous):

myininaya will who you how, because last time i did this i got in a huge discussion mmmm pie

myininaya (myininaya):

apple pie?

OpenStudy (anonymous):

o.o

myininaya (myininaya):

chicken pot pie

OpenStudy (anonymous):

ok fine. lets to it the easy (for me) way \[\frac{-337}{4}=-84\tfrac{1}{4}\] and since 2 divides 84 evenly this angle is co terminal with \[-\frac{\pi}{4}\]

OpenStudy (anonymous):

rhubarb pie

myininaya (myininaya):

what value does pie hold?

myininaya (myininaya):

does pie=pi?

OpenStudy (anonymous):

\[\sin(-\frac{\pi}{4})=-\frac{\sqrt{2}}{2}, \cos(-\frac{\pi}{4})=\frac{\sqrt{2}}{2}\] and since tangent is the ratio of sine over cosine you get -1

OpenStudy (anonymous):

so the final answer is -1?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

and if you doubt me, here is proof http://www.wolframalpha.com/input/?i=tan%28-337+pi%2F4%29

OpenStudy (anonymous):

\[\color{blue}{\text{sk}}\color{red}{\pi}\color{blue}{\text{y}}\]

OpenStudy (anonymous):

that, my myininaya, was for you

myininaya (myininaya):

pie?

OpenStudy (anonymous):

c'mon. what did i write?

myininaya (myininaya):

pi?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

what do you see there?

myininaya (myininaya):

lol this pie concerns me

myininaya (myininaya):

not your pi his pie

OpenStudy (anonymous):

i made a joke. visual pun. ok not that clever but still...